What is the % by volume (v/v) of solution if 7.5 L HCl is added to 1900 mL of water?


1
Expert's answer
2021-10-03T20:05:02-0400

We can use the volume of water and HCl solution expressed in liters to find the percentage in volume with respect to HCl:


"\%(v/v)=\cfrac{v_{HCl}}{v_{T}} \times 100
\\ \%(v/v) =\cfrac{v_{HCl}}{v_{HCl}+v_{water}}\times 100
\\ \%(v/v)=\cfrac{7.5\,L}{(7.5+1.9)\,L}\times 100 
\\ \%(v/v)=\cfrac{7.5\,\cancel{L}}{9.4\, \cancel{L} }\times 100
\\ \implies \%(v/v)=79.787\%\approxeq79.8\%"


In conclusion, the % by volume (v/v) for HCl is 79.8 %.

Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.